OpenStudy (anonymous):

A cube has an edge length of 2x. Write an expression for the total area of the cube.

OpenStudy (anonymous):

@Ddaddysgirl14

OpenStudy (anonymous):

well since all the edge lenghts are the same you would do 2x*2x to find one side then since there are 6 sides youd do (2x*2x)6

OpenStudy (anonymous):

i got 24x(cubed) @DaBest21

OpenStudy (anonymous):

well u have to distribute into the parentheses

OpenStudy (anonymous):

No, the answer i got was 24xcubed

OpenStudy (anonymous):

 - 2*4=8+15=24